% Generated by IEEEtran.bst, version: 1.14 (2015/08/26)
\begin{thebibliography}{10}
\providecommand{\url}[1]{#1}
\csname url@samestyle\endcsname
\providecommand{\newblock}{\relax}
\providecommand{\bibinfo}[2]{#2}
\providecommand{\BIBentrySTDinterwordspacing}{\spaceskip=0pt\relax}
\providecommand{\BIBentryALTinterwordstretchfactor}{4}
\providecommand{\BIBentryALTinterwordspacing}{\spaceskip=\fontdimen2\font plus
\BIBentryALTinterwordstretchfactor\fontdimen3\font minus
  \fontdimen4\font\relax}
\providecommand{\BIBforeignlanguage}[2]{{%
\expandafter\ifx\csname l@#1\endcsname\relax
\typeout{** WARNING: IEEEtran.bst: No hyphenation pattern has been}%
\typeout{** loaded for the language `#1'. Using the pattern for}%
\typeout{** the default language instead.}%
\else
\language=\csname l@#1\endcsname
\fi
#2}}
\providecommand{\BIBdecl}{\relax}
\BIBdecl

\bibitem{horodecki2009quantum}
R.~Horodecki, P.~Horodecki, M.~Horodecki, and K.~Horodecki, ``Quantum
  entanglement,'' \emph{Reviews of modern physics}, vol.~81, no.~2, p. 865,
  2009.

\bibitem{plenio2014introduction}
M.~B. Plenio and S.~S. Virmani, ``An introduction to entanglement theory,'' in
  \emph{Quantum Inf. Comput.}\hskip 1em plus 0.5em minus 0.4em\relax Springer,
  2014, pp. 173--209.

\bibitem{ekert1991quantum}
A.~K. Ekert, ``Quantum cryptography based on bell’s theorem,'' \emph{Physical
  review letters}, vol.~67, no.~6, p. 661, 1991.

\bibitem{bennett1993teleporting}
C.~H. Bennett, G.~Brassard, C.~Cr{\'e}peau, R.~Jozsa, A.~Peres, and W.~K.
  Wootters, ``Teleporting an unknown quantum state via dual classical and
  einstein-podolsky-rosen channels,'' \emph{Physical review letters}, vol.~70,
  no.~13, p. 1895, 1993.

\bibitem{bennett1992communication}
C.~H. Bennett and S.~J. Wiesner, ``Communication via one-and two-particle
  operators on einstein-podolsky-rosen states,'' \emph{Physical review
  letters}, vol.~69, no.~20, p. 2881, 1992.

\bibitem{lewenstein2001characterization}
M.~Lewenstein, B.~Kraus, P.~Horodecki, and J.~Cirac, ``Characterization of
  separable states and entanglement witnesses,'' \emph{Physical Review A},
  vol.~63, no.~4, p. 044304, 2001.

\bibitem{chen2003matrix}
K.~Chen and L.-A. Wu, ``A matrix realignment method for recognizing
  entanglement,'' \emph{Quantum Inf. Comput.}, no.~3, p. 193, 2003.

\bibitem{rudolph2005further}
O.~Rudolph, ``Further results on the cross norm criterion for separability,''
  \emph{Quantum Information Processing}, vol.~4, no.~3, pp. 219--239, 2005.

\bibitem{guhne2006entanglement}
O.~G{\"u}hne, M.~Mechler, G.~T{\'o}th, and P.~Adam, ``Entanglement criteria
  based on local uncertainty relations are strictly stronger than the
  computable cross norm criterion,'' \emph{Physical Review A}, vol.~74, no.~1,
  p. 010301, 2006.

\bibitem{zhang2008entanglement}
C.-J. Zhang, Y.-S. Zhang, S.~Zhang, and G.-C. Guo, ``Entanglement detection
  beyond the computable cross-norm or realignment criterion,'' \emph{Physical
  Review A}, vol.~77, no.~6, p. 060301, 2008.

\bibitem{spengler2012entanglement}
C.~Spengler, M.~Huber, S.~Brierley, T.~Adaktylos, and B.~C. Hiesmayr,
  ``Entanglement detection via mutually unbiased bases,'' \emph{Physical Review
  A}, vol.~86, no.~2, p. 022311, 2012.

\bibitem{shen2015separability}
S.-Q. Shen, M.-Y. Wang, M.~Li, and S.-M. Fei, ``Separability criteria based on
  the realignment of density matrices and reduced density matrices,''
  \emph{Physical Review A}, vol.~92, no.~4, p. 042332, 2015.

\bibitem{shang2018enhanced}
J.~Shang, A.~Asadian, H.~Zhu, and O.~G{\"u}hne, ``Enhanced entanglement
  criterion via informationally complete measurements,'' \emph{Physical Review
  A}, vol.~98, no.~2, p. 022309, 2018.

\bibitem{sarbicki2020family}
G.~Sarbicki, G.~Scala, and D.~Chru{\'s}ci{\'n}ski, ``Family of multipartite
  separability criteria based on a correlation tensor,'' \emph{Physical Review
  A}, vol. 101, no.~1, p. 012341, 2020.

\bibitem{shi2023family}
X.~Shi and Y.~Sun, ``A family of separability criteria and lower bounds of
  concurrence,'' \emph{Quantum Information Processing}, vol.~22, no.~3, p. 131,
  2023.

\bibitem{shi2024entanglement}
S.~Xian, ``The entanglement criteria based on equiangular tight frames,''
  \emph{Journal of Physics A: Mathematical and Theoretical}, vol.~57, no.~07,
  p. 075302, 2024.

\bibitem{wootters1998entanglement}
W.~K. Wootters, ``Entanglement of formation of an arbitrary state of two
  qubits,'' \emph{Physical Review Letters}, vol.~80, no.~10, p. 2245, 1998.

\bibitem{christandl2004squashed}
M.~Christandl and A.~Winter, ``“squashed entanglement”: an additive
  entanglement measure,'' \emph{Journal of mathematical physics}, vol.~45,
  no.~3, pp. 829--840, 2004.

\bibitem{chen2005concurrence}
K.~Chen, S.~Albeverio, and S.-M. Fei, ``Concurrence of arbitrary dimensional
  bipartite quantum states,'' \emph{Physical review letters}, vol.~95, no.~4,
  p. 040504, 2005.

\bibitem{de2007lower}
J.~I. de~Vicente, ``Lower bounds on concurrence and separability conditions,''
  \emph{Physical Review A}, vol.~75, no.~5, p. 052320, 2007.

\bibitem{li2020improved}
M.~Li, Z.~Wang, J.~Wang, S.~Shen, and S.-m. Fei, ``Improved lower bounds of
  concurrence and convex-roof extended negativity based on bloch
  representations,'' \emph{Quantum Information Processing}, vol.~19, no.~4, pp.
  1--11, 2020.

\bibitem{guhne2009entanglement}
O.~G{\"u}hne and G.~T{\'o}th, ``Entanglement detection,'' \emph{Physics
  Reports}, vol. 474, no. 1-6, pp. 1--75, 2009.

\bibitem{chruscinski2014entanglement}
D.~Chru{\'s}ci{\'n}ski and G.~Sarbicki, ``Entanglement witnesses: construction,
  analysis and classification,'' \emph{Journal of Physics A: Mathematical and
  Theoretical}, vol.~47, no.~48, p. 483001, 2014.

\bibitem{chruscinski2018entanglement}
D.~Chru{\'s}ci{\'n}ski, G.~Sarbicki, and F.~Wudarski, ``Entanglement witnesses
  from mutually unbiased bases,'' \emph{Physical Review A}, vol.~97, no.~3, p.
  032318, 2018.

\bibitem{li2019mutually}
T.~Li, L.-M. Lai, S.-M. Fei, and Z.-X. Wang, ``Mutually unbiased measurement
  based entanglement witnesses,'' \emph{International Journal of Theoretical
  Physics}, vol.~58, pp. 3973--3985, 2019.

\bibitem{siudzinska2021entanglement}
K.~Siudzi{\'n}ska and D.~Chru{\'s}ci{\'n}ski, ``Entanglement witnesses from
  mutually unbiased measurements,'' \emph{Scientific Reports}, vol.~11, no.~1,
  p. 22988, 2021.

\bibitem{siudzinska2022indecomposability}
K.~Siudzi{\'n}ska, ``Indecomposability of entanglement witnesses constructed
  from symmetric measurements,'' \emph{Scientific Reports}, vol.~12, no.~1, p.
  10785, 2022.

\bibitem{terhal2000schmidt}
B.~M. Terhal and P.~Horodecki, ``Schmidt number for density matrices,''
  \emph{Physical Review A}, vol.~61, no.~4, p. 040301, 2000.

\bibitem{cozzolino2019high}
D.~Cozzolino, B.~Da~Lio, D.~Bacco, and L.~K. Oxenl{\'o}we, ``High-dimensional
  quantum communication: benefits, progress, and future challenges,''
  \emph{Advanced Quantum Technologies}, vol.~2, no.~12, p. 1900038, 2019.

\bibitem{kues2017chip}
M.~Kues, C.~Reimer, P.~Roztocki, L.~R. Cort{\'e}s, S.~Sciara, B.~Wetzel,
  Y.~Zhang, A.~Cino, S.~T. Chu, B.~E. Little \emph{et~al.}, ``On-chip
  generation of high-dimensional entangled quantum states and their coherent
  control,'' \emph{Nature}, vol. 546, no. 7660, pp. 622--626, 2017.

\bibitem{wang2020qudits}
Y.~Wang, Z.~Hu, B.~C. Sanders, and S.~Kais, ``Qudits and high-dimensional
  quantum computing,'' \emph{Frontiers in Physics}, vol.~8, p. 589504, 2020.

\bibitem{paesani2021scheme}
S.~Paesani, J.~F. Bulmer, A.~E. Jones, R.~Santagati, and A.~Laing, ``Scheme for
  universal high-dimensional quantum computation with linear optics,''
  \emph{Physical Review Letters}, vol. 126, no.~23, p. 230504, 2021.

\bibitem{bavaresco2018measurements}
J.~Bavaresco, N.~Herrera~Valencia, C.~Kl{\"o}ckl, M.~Pivoluska, P.~Erker,
  N.~Friis, M.~Malik, and M.~Huber, ``Measurements in two bases are sufficient
  for certifying high-dimensional entanglement,'' \emph{Nature Physics},
  vol.~14, no.~10, pp. 1032--1037, 2018.

\bibitem{wyderka2023construction}
N.~Wyderka, G.~Chesi, H.~Kampermann, C.~Macchiavello, and D.~Bru{\ss},
  ``Construction of efficient schmidt-number witnesses for high-dimensional
  quantum states,'' \emph{Physical Review A}, vol. 107, no.~2, p. 022431, 2023.

\bibitem{liu2023characterizing}
S.~Liu, Q.~He, M.~Huber, O.~G{\"u}hne, and G.~Vitagliano, ``Characterizing
  entanglement dimensionality from randomized measurements,'' \emph{PRX
  Quantum}, vol.~4, no.~2, p. 020324, 2023.

\bibitem{simon2023}
S.~Morelli, M.~Huber, and A.~Tavakoli, ``Resource-efficient high-dimensional
  entanglement detection via symmetric projections,'' \emph{Physical Review
  Letters}, vol. 131, no.~17, p. 170201, 2023.

\bibitem{liu2024bounding}
S.~Liu, M.~Fadel, Q.~He, M.~Huber, and G.~Vitagliano, ``Bounding entanglement
  dimensionality from the covariance matrix,'' \emph{Quantum}, vol.~8, p. 1236,
  2024.

\bibitem{tavakoli2024enhanced}
A.~Tavakoli and S.~Morelli, ``Enhanced schmidt number criteria based on
  correlation trace norms,'' \emph{arXiv preprint arXiv:2402.09972}, 2024.

\bibitem{sanpera2001schmidt}
A.~Sanpera, D.~Bru{\ss}, and M.~Lewenstein, ``Schmidt-number witnesses and
  bound entanglement,'' \emph{Physical Review A}, vol.~63, no.~5, p. 050301,
  2001.

\bibitem{tomiyama1985geometry}
J.~Tomiyama, ``On the geometry of positive maps in matrix algebras. ii,''
  \emph{Linear algebra and its applications}, vol.~69, pp. 169--177, 1985.

\bibitem{horodecki2022five}
P.~Horodecki, L.~Rudnicki, and K.~{Z}yczkowski, ``Five open problems in quantum
  information theory,'' \emph{PRX Quantum}, vol.~3, no.~1, p. 010101, 2022.

\bibitem{zauner1999quantum}
G.~Zauner, ``Quantum designs,'' Ph.D. dissertation, Ph. D. thesis, University
  of Vienna Vienna, 1999.

\bibitem{scott2010symmetric}
A.~J. Scott and M.~Grassl, ``Symmetric informationally complete
  positive-operator-valued measures: A new computer study,'' \emph{Journal of
  Mathematical Physics}, vol.~51, no.~4, 2010.

\bibitem{scott2017sics}
A.~J. Scott, ``Sics: Extending the list of solutions,'' \emph{arXiv preprint
  arXiv:1703.03993}, 2017.

\bibitem{rastegin2014notes}
A.~E. Rastegin, ``Notes on general sic-povms,'' \emph{Physica Scripta},
  vol.~89, no.~8, p. 085101, 2014.

\bibitem{be}
I.~Bengtsson and K.~{\''Z}yczkowski, \emph{Geometry of quantum states: an
  introduction to quantum entanglement}, 2017.

\bibitem{wu2009entropic}
S.~Wu, S.~Yu, K.~M{\'o}lmer \emph{et~al.}, ``Entropic uncertainty relation for
  mutually unbiased bases,'' \emph{Physical Review A}, vol.~79, no.~2, p.
  022104, 2009.

\bibitem{shi2023lower}
S.~Xian, ``Lower bounds of entanglement quantifiers based on entanglement
  witnesses,'' \emph{arXiv preprint arXiv:2312.17620}, 2024.

\end{thebibliography}

\documentclass[pra,aps,showpacs,twocolumn,superscriptaddress]{revtex4}

\usepackage{amsmath,amsfonts,amssymb,caption,hyperref,color,epsfig,graphics,subfigure,graphicx,latexsym,mathrsfs,revsymb,theorem,url,verbatim,epstopdf}
\allowdisplaybreaks[3]

\hypersetup{colorlinks,linkcolor={blue},citecolor={blue},urlcolor={red}}

\usepackage{hyperref}

\makeatletter

\newcommand{\Rmnum}[1]{\expandafter\@slowromancap\romannumeral #1@}
\makeatother
\newtheorem{definition}{Definition}
\newtheorem{proposition}[definition]{Proposition}
\newtheorem{Lemma}[definition]{Lemma}

\newtheorem{Theorem}[definition]{Theorem}

\newtheorem{conjecture}[definition]{Conjecture}

\newtheorem{remark}[definition]{Remark}

\newtheorem{example}{Example}
\newtheorem{question}[definition]{Question}

\def\squareforqed{\hbox{\rlap{$\sqcap$}$\sqcup$}}
\def\qed{\ifmmode\squareforqed\else{\unskip\nobreak\hfil
		\penalty50\hskip1em\null\nobreak\hfil\squareforqed
		\parfillskip=0pt\finalhyphendemerits=0\endgraf}\fi}
\def\endenv{\ifmmode\;\else{\unskip\nobreak\hfil
		\penalty50\hskip1em\null\nobreak\hfil\;
		\parfillskip=0pt\finalhyphendemerits=0\endgraf}\fi}
\newenvironment{proof}{\noindent \textbf{{Proof.~} }}{\qed}
\def\Dbar{\leavevmode\lower.6ex\hbox to 0pt
	{\hskip-.23ex\accent"16\hss}D}
\makeatletter
\def\url@leostyle{%
	\@ifundefined{selectfont}{\def\UrlFont{\sf}}{\def\UrlFont{\small\ttfamily}}}
\makeatother
\def\bcj{\begin{conjecture}}
	\def\ecj{\end{conjecture}}
\def\bcr{\begin{corollary}}
	\def\ecr{\end{corollary}}
\def\bd{\begin{definition}}
	\def\ed{\end{definition}}
\def\bea{\begin{eqnarray}}
\def\eea{\end{eqnarray}}
\def\bem{\begin{enumerate}}
	\def\eem{\end{enumerate}}
\def\bex{\begin{example}}
	\def\eex{\end{example}}
\def\bim{\begin{itemize}}
	\def\eim{\end{itemize}}
\def\bl{\begin{lemma}}
	\def\el{\end{lemma}}
\def\bma{\begin{bmatrix}}
	\def\ema{\end{bmatrix}}
\def\bpf{\begin{proof}}
	\def\epf{\end{proof}}
\def\bpp{\begin{proposition}}
	\def\epp{\end{proposition}}
\def\bqu{\begin{question}}
	\def\equ{\end{question}}
\def\br{\begin{remark}}
	\def\er{\end{remark}}
\def\bt{\begin{theorem}}
	\def\et{\end{theorem}}

\def\btb{\begin{tabular}}
	\def\etb{\end{tabular}}

\newcommand{\nc}{\newcommand}

\def\a{\alpha}
\def\b{\beta}

\def\r{\rho}
\def\s{\sigma}

\def\o{\omega}

\nc{\bbA}{\mathbb{A}} \nc{\bbB}{\mathbb{B}} \nc{\bbC}{\mathbb{C}}
\nc{\bbD}{\mathbb{D}} \nc{\bbE}{\mathbb{E}} \nc{\bbF}{\mathbb{F}}
\nc{\bbG}{\mathbb{G}} \nc{\bbH}{\mathbb{H}} \nc{\bbI}{\mathbb{I}}
\nc{\bbJ}{\mathbb{J}} \nc{\bbK}{\mathbb{K}} \nc{\bbL}{\mathbb{L}}
\nc{\bbM}{\mathbb{M}} \nc{\bbN}{\mathbb{N}} \nc{\bbO}{\mathbb{O}}
\nc{\bbP}{\mathbb{P}} \nc{\bbQ}{\mathbb{Q}} \nc{\bbR}{\mathbb{R}}
\nc{\bbS}{\mathbb{S}} \nc{\bbT}{\mathbb{T}} \nc{\bbU}{\mathbb{U}}
\nc{\bbV}{\mathbb{V}} \nc{\bbW}{\mathbb{W}} \nc{\bbX}{\mathbb{X}}
\nc{\bbZ}{\mathbb{Z}}

\nc{\bA}{{\bf A}} \nc{\bB}{{\bf B}} \nc{\bC}{{\bf C}}
\nc{\bD}{{\bf D}} \nc{\bE}{{\bf E}} \nc{\bF}{{\bf F}}
\nc{\bG}{{\bf G}} \nc{\bH}{{\bf H}} \nc{\bI}{{\bf I}}
\nc{\bJ}{{\bf J}} \nc{\bK}{{\bf K}} \nc{\bL}{{\bf L}}
\nc{\bM}{{\bf M}} \nc{\bN}{{\bf N}} \nc{\bO}{{\bf O}}
\nc{\bP}{{\bf P}} \nc{\bQ}{{\bf Q}} \nc{\bR}{{\bf R}}
\nc{\bS}{{\bf S}} \nc{\bT}{{\bf T}} \nc{\bU}{{\bf U}}
\nc{\bV}{{\bf V}} \nc{\bW}{{\bf W}} \nc{\bX}{{\bf X}}
\nc{\bZ}{{\bf Z}}

\nc{\cA}{{\cal A}} \nc{\cB}{{\cal B}} \nc{\cC}{{\cal C}}
\nc{\cD}{{\cal D}} \nc{\cE}{{\cal E}} \nc{\cF}{{\cal F}}
\nc{\cG}{{\cal G}} \nc{\cH}{{\cal H}} \nc{\cI}{{\cal I}}
\nc{\cJ}{{\cal J}} \nc{\cK}{{\cal K}} \nc{\cL}{{\cal L}}
\nc{\cM}{{\cal M}} \nc{\cN}{{\cal N}} \nc{\cO}{{\cal O}}
\nc{\cP}{{\cal P}} \nc{\cQ}{{\cal Q}} \nc{\cR}{{\cal R}}
\nc{\cS}{{\cal S}} \nc{\cT}{{\cal T}} \nc{\cU}{{\cal U}}
\nc{\cV}{{\cal V}} \nc{\cW}{{\cal W}} \nc{\cX}{{\cal X}}
\nc{\cZ}{{\cal Z}}

\nc{\hA}{{\hat{A}}} \nc{\hB}{{\hat{B}}} \nc{\hC}{{\hat{C}}}
\nc{\hD}{{\hat{D}}} \nc{\hE}{{\hat{E}}} \nc{\hF}{{\hat{F}}}
\nc{\hG}{{\hat{G}}} \nc{\hH}{{\hat{H}}} \nc{\hI}{{\hat{I}}}
\nc{\hJ}{{\hat{J}}} \nc{\hK}{{\hat{K}}} \nc{\hL}{{\hat{L}}}
\nc{\hM}{{\hat{M}}} \nc{\hN}{{\hat{N}}} \nc{\hO}{{\hat{O}}}
\nc{\hP}{{\hat{P}}} \nc{\hR}{{\hat{R}}} \nc{\hS}{{\hat{S}}}
\nc{\hT}{{\hat{T}}} \nc{\hU}{{\hat{U}}} \nc{\hV}{{\hat{V}}}
\nc{\hW}{{\hat{W}}} \nc{\hX}{{\hat{X}}} \nc{\hZ}{{\hat{Z}}}

\nc{\hn}{{\hat{n}}}

\def\max{\mathop{\rm max}}
\def\min{\mathop{\rm min}}

\def\tr{\mathop{\rm Tr}}

\newcommand{\bra}[1]{\langle#1|}
\newcommand{\ket}[1]{|#1\rangle}

\newcommand{\norm}[1]{\lVert#1\rVert}

\def\Dbar{\leavevmode\lower.6ex\hbox to 0pt
	{\hskip-.23ex\accent"16\hss}D}

\begin{document}
	\title{Families of Schmidt-number witnesses for high dimensional quantum states}
	
	\author{Xian Shi}\email[]
	{shixian01@gmail.com}
	\affiliation{College of Information Science and Technology,
		Beijing University of Chemical Technology, Beijing 100029, China}

	%
	
	
	
	\date{\today}
	\begin{abstract}
		Higher dimensional entangled states demonstrate significant advantages in quantum information processing tasks. Schmidt number is a quantity on the entanglement dimension of a bipartite state. Here we build families of $k$-positive maps from the symmetric information complete positive operator-valued measurements and mutually unbiased bases, and we also present the Schmidt number witnesses, correspondingly. At last, based on the witnesses obtained from mutually unbiased bases, we show the distance between a bipartite state and the set of states with Schmidt number less than $k$.
	\end{abstract}

	\pacs{03.65.Ud, 03.67.Mn}
	\maketitle

\section{Introduction}
\indent	Entanglement is one of the most fundamental features in quantum mechanics compared to classical physics \cite{horodecki2009quantum,plenio2014introduction}. It also plays critical roles in quantum information and quantum computation theory, such as quantum cryptography \cite{ekert1991quantum}, teleportation \cite{bennett1993teleporting}, and superdense coding \cite{bennett1992communication}. \par
From the start of quantum information theory, lots of efforts have been devoted to the problems of distinguishing whether a state is separable or entangled \cite{lewenstein2001characterization,chen2003matrix,rudolph2005further,guhne2006entanglement,zhang2008entanglement,spengler2012entanglement,shen2015separability,shang2018enhanced,sarbicki2020family,shi2023family,shi2024entanglement} and quantifying entanglement of the state \cite{wootters1998entanglement,christandl2004squashed,chen2005concurrence,de2007lower,li2020improved}. A commonly used method to certify the entanglement of a state is to build an effective entanglement witness. There are many ways to construct the entanglement witness \cite{guhne2009entanglement,chruscinski2014entanglement}. In 2018, Chruscinski $et$ $al.$ showed a method to construct entanglement witnesses from mutually unbiased bases (MUB) \cite{chruscinski2018entanglement}. Whereafter, the authors in \cite{li2019mutually,siudzinska2021entanglement,siudzinska2022indecomposability} generalized the method to build entanglement witnesses with other classes of positive operator-valued measurements (POVM).  One of the most essential entanglement measures is the Schmidt number (SN) \cite{terhal2000schmidt}, this quantity indicates the lowest dimension of the system needed 
to generate the entanglement. Furthermore, genuine high dimensional entanglement plays important roles in many quantum information tasks, such as, quantum communication \cite{cozzolino2019high}, quantum control \cite{kues2017chip} and universal quantum computation \cite{wang2020qudits,paesani2021scheme}. 

However, like most entanglement measures, it is hard to obtain the SN of a generic entangled state. Recently, the method to bound the SN of an entangled state attracted much attention from the relevant researchers \cite{bavaresco2018measurements,wyderka2023construction,liu2023characterizing,simon2023,liu2024bounding,tavakoli2024enhanced}. In \cite{bavaresco2018measurements}, Bavaresco $et$ $al.$ proposed a method to bound the dimension of an entangled state. Recently, Liu $et$ $al.$ presented the results of SN of a given state based on its covariance matrix \cite{liu2023characterizing,liu2024bounding}. Tavakoli and Morelli showed the bound of SN of a given state with the help of MUBs and SIC POVMs \cite{tavakoli2024enhanced}. Similar to entanglement, a straightford method to certify the dimension of an entangled state is by constructing the $k$-postive maps \cite{terhal2000schmidt} or SN witnesses \cite{terhal2000schmidt,sanpera2001schmidt,wyderka2023construction}. However, there are few results obtained on constructing the SN witnesses with the help of certain POVMs.

In this manuscript, we will present the methods to construct $k$-positive maps with the use of symmetric information complete (SIC) POVMs and MUBs, which generalizes the methods of \cite{chruscinski2018entanglement}. we also give the corresponding SN witness. Moreover, we compare the $k$-positive maps here with those in \cite{terhal2000schmidt,tomiyama1985geometry}. At last, we present the lower bounds of the distance between a bipartite mixed state and the set of states with Schmidt number less than $k$ based on the SN witnesses, which is built from MUBs. 
\section{Preliminary Knowledge}
In the manuscript, the quantum systems we considered here are finite dimensions. Next we denote $\mathcal{D}(\mathcal{H}_{AB})$ as the set consisting of the states of $\mathcal{H}_{AB},$  $$\mathcal{D}(\mathcal{H}_{AB})=\{\rho_{AB}|\rho_{AB}\ge0,\tr\rho_{AB}=1\}.$$ And we denote $\ket{\psi_d}=\sum_{i=0}^{d-1}\ket{ii}$ as the maximally entangled states of $\mathcal{H}_{AB}$ with $\mathrm{Dim}(\mathcal{H}_A)=\mathrm{Dim}(\mathcal{H}_B)=d.$

In this section, we will first recall the knowledge of the Schmidt number for a bipartite mixed state, then we will recall the definition and properties of SICs and MUBs, correspondingly. At last, we will present the definition of the distance to the set of states with Schmidt number less than k, $D_k(\cdot).$
\subsection{Schmidt Number}
Assume $\ket{\psi}_{AB}=\sum_{ij}c_{ij}\ket{ij}$ is a pure state in $\mathcal{H}_{AB}$ with $\mathrm{Dim}(\mathcal{H}_A)=d_A$ and $\mathrm{Dim}(\mathcal{H}_B)=d_B.$ 
There always exists orthonormal bases $\{\ket{\tilde{i}}_A\}$ and $\{\ket{\tilde{i}}_B\}$ in $\mathcal{H}_A$ and $\mathcal{H}_B,$ respectively such that $$\ket{\psi}_{AB}=\sum_{i=1}^k\sqrt{\lambda_i}\ket{\tilde{i}\tilde{i}},$$ here $\lambda_i>0$ and $\sum_i\lambda_i^2=1.$ Here the nonzero number $k$ is called the Schmidt number of $\ket{\psi}$ \cite{terhal2000schmidt}, $i.$ $e.$, $SN(\ket{\psi})=k.$ The Schmidt number of a mixed state $\rho_{AB}$ is defined as follows \cite{terhal2000schmidt},
\begin{align}
SN(\rho)=\min_{\rho=\sum_ip_i\ket{\psi_i}\bra{\psi_i}}\max_i SR(\ket{\psi_i}),
\end{align}
where the minimization takes over all the decompositions of $\rho_{AB}=\sum_ip_i\ket{\psi_i}\bra{\psi_i}.$ The Schmidt number is entanglement monotone, and it can be seen as a key quantity on the power of entanglement resources.\par 
Through the definition of Schmidt number, one can classify the states of $\mathcal{H}_{AB}$ as follows. Let 
\begin{align}
S_k=\{\rho|SN(\rho)\le k\},
\end{align}
due to the definition of $S_k$, we have $S_k\subset S_{k+1},$ and $S_1$ is the set of separable states.  Assume $\rho_{AB}\in S_k$ is a bipartite state in $\mathcal{H}_{AB}$ with $\mathrm{Dim}(\mathcal{H}_A)=\mathrm{Dim}(\mathcal{H}_B)=d$, the authors in \cite{terhal2000schmidt} showed that
\begin{align}
\tr(\rho\ket{\psi_d}\bra{\psi_d})\le \frac{k}{d},\label{f1}
\end{align}
 Besides, from the definition of $S_k$, we have $S_k$ is a convex set. Hence, we could construct the Schmidt number witness $W_k$ to validate a bipartite state $\rho_{AB}\in S_{k+1},$ due to the Hahn-Banach theorem, if
\begin{align*}
\tr(W_k\rho_k)\ge& 0 \hspace{5mm} \forall \rho_k\in S_{k},\\
\tr(W_k\rho)<& 0\hspace{5mm} \exists \rho\in\mathcal{D}(\mathcal{H}_{AB}),
\end{align*}
then we call $W_k$ a SN-(k+1) witness.

Then we recall the following result on the Schmidt number of a bipartite state obtained in \cite{terhal2000schmidt},
\begin{Lemma}\label{l1}
	Assume $\rho$ is a bipartite state on $\mathcal{H}_{AB}.$ $\rho$ has Schmidt number at least $k+1$ if and only if there exists a $k$-positive linear map $\Lambda_k$ such that 
	\begin{align}
	(I\otimes\Lambda_k)(\rho)\ngeq0.\label{kp}
	\end{align}
	The linear Hermitticity-preserving map $\Lambda$ is $k$-positive if and only if 
	\begin{align}
	(I\otimes\Lambda)(\ket{\psi_k}\bra{\psi_k})\ge0,
	\end{align}
	here $\ket{\psi_k}$ are arbitrary maximally entangled state with Schmidt number $k.$ Finally, if $\Lambda$ is $k$-positive, then $\Lambda^{\dagger}$ defined by $\tr A^{\dagger}\Lambda(B)=\tr\Lambda^{\dagger}(A^{\dagger})B$ for all $A$ and $B$, is also $k$-positive.
\end{Lemma}

In \cite{tomiyama1985geometry,terhal2000schmidt}, the authors showed a family of positive maps $\Lambda_p(X)$ with the following form
\begin{align}
\Lambda_p(X)=\tr(X)\mathbb{I}-pX,\label{klp}
\end{align}
where $X$ is a linear operator of $\mathcal{H}$. When $k\ge \frac{1}{p}>k+1$, $\Lambda_p$ is $k$-positive.
\subsection{SICs and MUBs}
Assume $\mathcal{H}_d$ is a Hilbert space with dimension $d,$ $\{{E_i}=\frac{1}{d}\ket{\phi_i}\bra{\phi_i}|i=1,2,\cdots,d^2\}$ is a positive operator valued measure(POVM) of $\mathcal{H}_d,$ here $\ket{\phi_i}$ are pure states with
\begin{align}
|\bra{\phi_j}\phi_k\rangle|^2=\frac{1}{d+1},\hspace{4mm}\forall j\ne k,
\end{align}
then $\{E_i\}_{i=1}^{d^2}$ is a SIC-POVM. The existence of SIC-POVMs in every dimension is still an open problem \cite{horodecki2022five}, readers who are interesting to the problem can refer to \cite{zauner1999quantum,scott2010symmetric,scott2017sics}. 

Assume $\rho$ is a state of $\mathcal{D}(\mathcal{H}_d),$ $\{E_i|i=1,2,\cdots,d^2\}$ is a SIC-POVM, then 
\begin{align}
\sum_{j=1}^{d^2}|\tr(P_j\rho)|^2=\frac{\tr\rho^2+1}{d+d^2},
\end{align}
which is showed in \cite{rastegin2014notes}.

Next we recall the definition of MUBs. Assume $\{\ket{e^l_i}|i=1,2,\cdots,d\}_{l=1}^L$ are $L$ orthonormal bases, and 
\begin{align*}
|\bra{e^m_j}f^{m^{'}}_k\rangle|^2=&\frac{1}{d}\hspace{5mm}\forall j,k\hspace{5mm} \textit{ m$\ne $$m^{'}$},\\
|\bra{e^m_j}f^{m}_k\rangle|^2=&\delta_{jk}.
\end{align*} 
then they are MUBs. For any space with dimension $d,$ there exists at most $d+1$ MUBs. If the upper bound is reached, the set of MUBs is called a complete set. It is well known that the complete sets of MUBs are existed when the dimension of the Hilbert space is a number with prime power. However, the existence of the complete sets of MUBs is unknown for arbitrary dimensional systems, even if the dimension is 6 \cite{horodecki2022five}.

Let $\{Q^{(\a)}_i=\ket{e^{(\a)}_i}\bra{e^{(\a)}_i}|i=1,2,\cdots,d\}_{\a=1}^L$ are $L$ MUBs and $\rho$ is a state in $\mathcal{D}(\mathcal{H}_d)$, then 
\begin{align}
	\sum_{\a=1}^L\sum_{i=1}^d|\tr(\r Q_i^{(\a)})|^2\le \tr(\r^2)+\frac{L-1}{d},
\end{align}
the above inequality is obtained in \cite{wu2009entropic}.

At last, we present the distance to the set $S_k$ for a bipartite state, $D_{k}(\cdot)$. Assume $\rho_{AB}$ is a bipartite mixed state, its distance to the set $S_k$ in terms of the Frobenius norm is defined as
\begin{align}
D_k(\rho)=\min_{\s\in S_k}\norm{\r-\s}_F,
\end{align}
where the minimum takes over all the state in $S_k.$
\section{k-positive maps based on SICs and MUBs}
Assume $\mathcal{H}_d$ is a Hilbert space with dimension $d,$ $\mathcal{M}=\{P_i=\frac{1}{d}\ket{\phi_i}\bra{\phi_i}\}_{i=1}^{d^2}$ is a SIC-POVMs. Next we present a class of $k$-positive maps $\Lambda(\cdot)$ with the help of the SIC-POVM $\mathcal{M}$. Let $\mathcal{O}$ be an orthogonal rotation in $\mathbb{R}^d$ around the axis $n_{*}=\frac{(1,1,\cdots,1)}{\sqrt{d}}$, that is, $\mathcal{O}n_{*}=n_{*},$
\begin{widetext}
\begin{align}
\Lambda(X)=  \frac{\mathbb{I}_d}{d}\tr(X)-h\sum_{g,l=1}^{d^2}\mathcal{O}_{gl}\tr[(X-\frac{\mathbb{I}_d}{d}\tr(X))P_l]P_g,\label{f2}
\end{align}
\end{widetext}
here $h=\sqrt{\frac{d^4+d^3}{(kd-1)(kd+k-2)}},$ $P_l$ and $P_k$ take over all the elements in a SIC-POVM $\mathcal{M}.$
\begin{Theorem}\label{t1}
	The map $\Lambda(\cdot)$ defined in (\ref{f2}) is $k$-positive.
\end{Theorem}
\begin{proof}
Due to the Lemma \ref{l1}, when we prove 
\begin{align}
(I\otimes \Lambda)(\ket{\psi_k}\bra{\psi_k})\ge 0,\label{f3}
\end{align}
for all maximally entangled state $\ket{\psi_k}$ with Schmidt number $k$, then we finish the proof. Here we utilize the following fact to prove (\ref{f3}): when $\rho$ is a Hermite matrix with trace 1, if $$\tr\rho^2\le \frac{1}{d-1},$$ then $\rho$ is a state \cite{be}.

Assume $\ket{\psi_k}=(U\otimes V)\sum_{i=0}^{k-1}\sqrt{\frac{1}{k}}\ket{ii}$, here $U$ and $V$ are arbitrary unitary operators of $\mathcal{H}_A$ and $\mathcal{H}_B$, respectively, then
\begin{align}
&\tr[(I\otimes \Lambda)(U\otimes V)(\ket{\psi_k}\bra{\psi_k})(U^{\dagger}\otimes V^{\dagger})]^2\nonumber\\
=&\tr\frac{1}{k^2}[(I\otimes\Lambda)(U\otimes V)(\sum_{i=0}^{k-1}\sum_{j=0}^{k-1}\ket{ii}\bra{jj})(U^{\dagger}\otimes V^{\dagger})]^2\nonumber\\
=&\tr\frac{1}{k^2}[\sum_{i,j=0}^{k-1}U\ket{i}\bra{j}U^{\dagger}\otimes\Lambda(V\ket{i}\bra{j}V^{\dagger})]^2\nonumber\\
=&\tr\frac{1}{k^2}[\sum_{i,j,m=0}^{k-1}U\ket{i}\bra{m}U^{\dagger}\otimes \Lambda(V\ket{i}\bra{j}V^{\dagger})\Lambda(V\ket{j}\bra{m}V^{\dagger})]\nonumber\\
=&\frac{1}{k^2}\tr\sum_{i,j=0}^{k-1}\Lambda(V\ket{i}\bra{j}V^{\dagger})\Lambda(V\ket{j}\bra{i}V^{\dagger})\label{f4}.
\end{align} 
Next we compute $\Lambda(\ket{i}\bra{j}),$ when $i=j$,
\begin{align}
\Lambda(V\ket{i}\bra{i}V^{\dagger})=\frac{\mathbb{I}_d}{d}-h\sum_{g,l=1}^{d^2}\mathcal{O}_{gl}\tr[(V\ket{i}\bra{i}V^{\dagger}-\frac{\mathbb{I}_d}{d})P_l]P_g,
\end{align}
when $i\ne j$,
\begin{align}
\Lambda(V\ket{i}\bra{j}V^{\dagger})=-h\sum_{g,l=1}^{d^2}\mathcal{O}_{gl}\tr[(V\ket{i}\bra{j}V^{\dagger})P_l]P_g.
\end{align}
\begin{widetext}
\begin{align}
(\ref{f4})=&\frac{1}{kd}+\frac{h^2}{k^2}\sum_{i=0}^{k-1}\sum_{g,l,s,t=1}^{d^2}\mathcal{O}_{gl}\mathcal{O}_{ts}\tr[(V\ket{i}\bra{i}V^{\dagger}-\frac{\mathbb{I}_d}{d})P_l]\tr[(V\ket{i}\bra{i}V^{\dagger}-\frac{\mathbb{I}_d}{d})P_s]\tr(P_gP_t)\nonumber\\
-&\frac{2h}{k^2}\sum_{i=0}^{k-1}\tr\sum_{g,l=1}^{d^2}\mathcal{O}_{gl}\tr[(V\ket{i}\bra{i}V^{\dagger}-\frac{\mathbb{I}_d}{d})P_l]\tr(P_g)+\frac{h^2}{k^2}\sum_{i\ne j}\sum_{g,l,s,t=1}^{d^2}\mathcal{O}_{gl}\mathcal{O}_{st}\tr[V\ket{i}\bra{j}V^{\dagger}P_l]\tr[V\ket{j}\bra{i}V^{\dagger}P_t]\tr(P_gP_s)\nonumber\\
=&\frac{1}{kd}+\frac{h^2}{d^2k^2}\sum_{i=0}^{k-1}\sum_{l=1}^{d^2}|\tr[(V\ket{i}\bra{i}V^{\dagger}-\frac{\mathbb{I}_d}{d})P_l]|^2+\frac{h^2}{k^2d^2}\sum_{i\ne j}\sum_{l=1}^{d^2}\tr(P_l V\ket{i}\bra{j}V^{\dagger})\tr(P_l V\ket{j}\bra{i}V^{\dagger})\nonumber\\
\le& \frac{1}{kd}+\frac{h^2(d-1)}{d^4(d+1)k}+\frac{h^2(k-1)}{d^4k}\nonumber\\
=&\frac{d^4+d^3+h^2kd+kh^2-2h^2}{kd^4(d+1)}\nonumber\\
=&\frac{1}{dk-1},
\end{align}
\end{widetext}
in the last inequality, we apply Lemma \ref{l2} in Sec. \ref{app} and the following derived in \cite{rastegin2014notes}
\begin{align}
\sum_{j=1}^{d^2}|\tr(P_j\rho)|^2=\frac{\tr\rho^2+1}{d+d^2},
\end{align}
hence we have $\Psi(\cdot)$ is $k$-positive.
\end{proof}

Based on Lemma \ref{l1}, we can provide a class of witnesses on detecting whether a bipartite state is in $S_k$ through the $k$-positive map defined in (\ref{f2}),
\begin{align}
W_k=\frac{h+d}{d^2}\mathbb{I}_d\otimes \mathbb{I}_d-h\sum_{gl}\mathcal{O}_{gl}\overline{P_l}\otimes P_g,\label{w1}
\end{align}
here $h=\sqrt{\frac{d^4+d^3}{(kd-1)(kd+k-2)}}.$

Next we present a class of $k$-positive maps $\Lambda(\cdot)$ based on the MUBs. Let $\{\ket{e^{\a}_i}|\a=1,2,\cdots,m\}_{i=1}^L$ be the MUBs, $\mathcal{O}^{(\a)}$ be $L$ orthogonal rotation in $\mathbb{R}^d$ around the axis $n_{*}=\frac{1}{d}(1,1,\cdots,1),$ that is, $\mathcal{O}^{(\a)}n_{*}=n_{*}.$ Based on the MUBs, we also present a set of $k$-positive maps,
\begin{widetext}
\begin{align}
\Theta_k(X)=  \frac{\mathbb{I}_d}{d}\tr(X)-h_s\sum_{\a=1}^L\sum_{g,l=1}^{d}\mathcal{O}^{(\a)}_{gl}\tr[(X-\frac{\mathbb{I}_d}{d}\tr(X))Q^{(\a)}_l]Q^{(\a)}_g,\label{f5}
\end{align}
\end{widetext}
here $h_s=\sqrt{\frac{1}{(dk-1)(Lk-L+d-1)}},$ $Q_l^{(\a)}=\ket{e_i^{\a}}\bra{e_i^{\a}}$, $\{\ket{e_i^{\a}}|\a=1,2,\cdots,L\}_{i=1}^d$ are $L$ MUBs. 
\begin{Theorem}
$\Theta_k(\cdot)$ defined in (\ref{f5}) are $k$-positive. 
\end{Theorem}
\begin{proof}
	Here we apply similar method of the proof of Theorem \ref{t1}. Assume $\ket{\psi_k}=(U\otimes V)\sum_{i=0}^{k-1}\sqrt{\frac{1}{k}}\ket{ii}$, here $U$ and $V$ are arbitrary unitary operators of $\mathcal{H}_A$ and $\mathcal{H}_B$, respectively, if $\tr[(I_k\otimes\Theta_k)(\ket{\psi_k}\bra{\psi_k})]^2\le \frac{1}{dk-1}$, then we finish the proof \cite{be}.
	\begin{align}
	&\tr[(I\otimes \Theta_k)(U\otimes V)(\ket{\psi_k}\bra{\psi_k})(U^{\dagger}\otimes V^{\dagger})]^2\nonumber\\
=&\frac{1}{k^2}\tr\sum_{i,j=0}^{k-1}\Theta_k(V\ket{i}\bra{j}V^{\dagger})\Theta_k(V\ket{j}\bra{i}V^{\dagger})\label{f6}
	\end{align}
	Next we compute $\Theta_k(V\ket{i}\bra{j}V^{\dagger}),$ when $i=j,$
	\begin{align}
	&\Theta_k(V\ket{i}\bra{i}V^{\dagger})\nonumber\\
	=&\frac{\mathbb{I}_d}{d}-h_s\sum_{\a=1}^L\sum_{g,l=1}^{d}\mathcal{O}^{(\a)}_{gl}\tr[(V\ket{i}\bra{i}V^{\dagger}-\frac{\mathbb{I}_d}{d})Q^{(\a)}_l]Q^{(\a)}_g,\label{f7}
	\end{align}
	when $i\ne j$,
	\begin{align}
	&\Theta_k(V\ket{i}\bra{j}V^{\dagger})\nonumber\\
	=&-h_s\sum_{\a=1}^L\sum_{g,l=1}^{d}\mathcal{O}^{(\a)}_{gl}\tr[(V\ket{i}\bra{j}V^{\dagger})Q^{(\a)}_l]Q^{(\a)}_g,\label{f8}
	\end{align}
	then based on (\ref{f7}) and (\ref{f8}), we have
	\begin{widetext}
	\begin{align*}
	(\ref{f6})\le &\frac{1}{k^2}[\frac{k}{d}+h_s^2\sum_{i=1}^k\sum_{\a=1}^L\sum_{n=1}^d|\tr(V\ket{i}\bra{i}V^{\dagger}-\frac{\mathbb{I}_d}{d})Q_n^{(\a)}|^2+h_s^2\frac{Lk(k-1)}{d}]\nonumber\\
	\le &\frac{1}{k^2}[\frac{k}{d}+kh_s^2(1-\frac{1}{d})+h_s^2\frac{Lk(k-1)}{d}]=\frac{1}{dk-1},
	\end{align*}
	In the inequalities, we have used Lemma \ref{l3} in Sec. \ref{app} the following derived in \cite{wu2009entropic}
	\begin{align}
	\sum_{\a=1}^L\sum_{l=1}^d|\tr(\r Q_l^{(\a)})|^2\le \tr(\r^2)+\frac{L-1}{d}
	\end{align}
	\end{widetext}
\end{proof}
Based on Lemma \ref{l1}, we can provide a class of witnesses on detecting whether a bipartite state is in $S_k$ through the $k$-positive map defined in (\ref{f5}),
\begin{align}
W_k=\frac{1+Lh_s}{d}\mathbb{I}_d\otimes \mathbb{I}_d-h_s\sum_{g,l=1}^d\sum_{\a=1}^L\mathcal{O}_{gl}^{(\a)}\overline{Q^{(\a)}_l}\otimes Q^{(\a)}_g,\label{w2}
\end{align}
here $h_s=\sqrt{\frac{1}{(dk-1)(Lk-L+d-1)}}.$
\begin{remark}
 When $k=1$, $h_s=\frac{1}{d-1},$ 
\begin{align}
W_k=\frac{d+L-1}{d(d-1)}\mathbb{I}_d\otimes \mathbb{I}_d-\frac{1}{d-1}\sum_{g,l=1}^d\sum_{\a=1}^L\mathcal{O}_{gl}^{(\a)}\overline{Q^{(\a)}_l}\otimes Q^{(\a)}_g,
\end{align}
which is the entanglement witnesses shown in \cite{chruscinski2018entanglement}. 
\end{remark}
\begin{remark}
	When $L=d+1,$ and $\mathcal{O}=\mathbb{I,}$ the $k$-positive map (\ref{f5}) can be written as
	\begin{align}
	W_k=\frac{\mathbb{I}_d}{d}{(1+h_c)}\tr(X)-h_cX,
	\end{align}
	here $h_c=\sqrt{\frac{1}{(dk-1)(kd+k-2)}}$, and 
	\begin{align*}
	\frac{1+h_c}{dh_c}=\frac{\frac{1}{h_c}+1}{d}=\frac{\sqrt{(dk-1)(kd+k-2)}+1}{d},
	\end{align*}
	as $\sqrt{(dk-1)(kd+k-2)}\in [k,k+1),$ we have this class of $k$-positive maps constructed from complete sets of MUBs are belong to the family of (\ref{klp}) in \cite{terhal2000schmidt}.
\end{remark}

\section{Applications}
\indent In this section, we will present the distance between the state and the set $S_k$ based on the witness obtained in the last section, the method here is based on \cite{shi2023lower}.

Assume $\rho$ is a bipartite mixed state, $Y_k$ is a Schmidt number $k$ witness of $\mathcal{H}_d\otimes\mathcal{H}_d$, let $a=\frac{\tr(Y_k)}{d^2}$, $b=\sqrt{\tr(Y_k^{\dagger}Y_k)-\frac{(\tr Y_k)^2}{d^2}},$ $V_k=\frac{Y_k-a\mathbb{I}\otimes\mathbb{I}}{b}$,
\begin{align}
D_{k}(\rho_{AB})=&\min_{\s\in S_{k}}\max_{\norm{W}_F=1}\tr(W(\rho_{AB}-\s_{AB}))\nonumber\\
\ge&|\tr[V_k(\rho-\omega)]\nonumber\\
=&|\tr[\frac{Y_k}{b}(\r-\o)-\frac{a}{b}(\r-\o)]\nonumber\\
=&|\tr[\frac{Y_k}{b}(\r-\o)]\nonumber\\
\ge&-\frac{1}{b}\tr Y_k\r,\label{wi}
\end{align}
in the first inequality, $\o$ is the optimal state in $S_k$, $V_k$ is a Hermite operator and $\norm{V_k}_F=1$. The last inequality is due to that $\o\in S_k,$ $\tr Y_k\o\ge0.$ 

Next we apply the witnesses obtained in (\ref{w2}) to show the distance between a bipartite state and the set $S_k$ relying on (\ref{wi}).
\begin{Theorem}
	Assume $\rho_{AB}$ is a bipartite mixed state on $\mathcal{H}_d\otimes\mathcal{H}_d$, let $\{\ket{e^{\a}_i}|\a=1,2,\cdots,m\}_{i=1}^L$ be the MUBs of the system $\mathcal{H}_d.$
Let $W_k$ be defined in (\ref{w2}), then
\begin{align}
D_k(\r)\ge -\frac{1}{\sqrt{h_s^2(Ld-L)}}\tr W_k\r
\end{align}
here $h_s=\sqrt{\frac{1}{(dk-1)(Lk-L+d-1)}}.$
\end{Theorem}
\begin{proof}
	Due to the (\ref{wi}), to obtain the lower bound of $D_k(\rho)$, we only need to compute the value of $b$. Let $Z_k=\sum_{g,l=1}^d\sum_{\a=1}^L\mathcal{O}_{gl}^{(\a)}\overline{Q^{(\a)}_l}\otimes Q^{(\a)}_g,$ 
	\begin{align}
&\tr Z_k\nonumber\\
=&\sum_{g,l=1}^d\sum_{\a=1}^L\mathcal{O}_{gl}^{(\a)}\tr\overline{Q^{(\a)}_l}\otimes Q^{(\a)}_g,\nonumber\\
=&\sum_{g,l=1}^d\sum_{\a=1}^L\mathcal{O}_{gl}^{(\a)}\nonumber\\
=&Ld,\label{y1}
\\
&\tr Z_k^{\dagger}Z_k\nonumber\\
=&\sum_{g,l,m,n=1}^d\sum_{\a,\b=1}^L\mathcal{O}_{gl}^{(\a)}\mathcal{O}_{mn}^{(\b)}\tr\overline{Q^{(\a)}_l}\overline{Q^{(\b)}_n}\otimes Q^{(\a)}_gQ^{(\b)}_m\nonumber\\
=&Ld+L^2-L,
\end{align}
then 
\begin{align}
b=&\sqrt{\tr(W_k^{\dagger}W_k)-\frac{(\tr W_k)^2}{d^2}}\nonumber\\
=&\sqrt{h_s^2(Ld-L)},
\end{align}
hence
\begin{align}
D_k(\r_{AB})\ge -\frac{1}{\sqrt{h_s^2(Ld-L)}}\tr W_k\r.
\end{align}
\end{proof}
\section{Conclusion}
Here we have presented families of $k$-positive maps in arbitrary dimensional systems based on the SIC POVMs and MUBs. Based on the $k$-positive maps, we also built the SN witnesses correspondingly. Then we compared the $k$-positive maps built from MUBs and the existing results. When $k=1,$ we found that the conclusion degrades into the map obtained in \cite{chruscinski2018entanglement}. When the $k$-positive maps built from a complete set of MUBs, they belong to the family obtained in \cite{terhal2000schmidt}. At last, we presented a defition of the distance between a bipartite state and the set $S_k.$ Moreover, we showed a lower bound of the distance based on the SN witnesses constructed from MUBs. Due to the important roles that higher dimensional systems played, our results can provide a reference for future work on the study of entanglement theory.

  \section{Acknowledgement}
X. S. was supported by the National Natural Science Foundation of China (Grant No. 12301580) and the Funds of College of Information Science and Technology, Beijing University of Chemical Technology (Grant No. 0104/11170044115).
\bibliographystyle{IEEEtran}
\bibliography{ref}

\section{Appendix}\label{app}
\begin{Lemma}\label{l2}
	Let $\mathcal{M}=\{M_j=\frac{1}{d}\ket{\phi_j}\bra{\phi_j}\}$ be a SIC-POVM in a $d$ dimensional system $\mathcal{H}$, and $\{\ket{i}|i=1,2,\cdots,d\}$ is any orthonormal base of $\mathcal{H}$, when $i\ne j,$
	\begin{align*}
	\sum_{j=1}^{d^2}\tr(M_j\ket{i}\bra{j})=\frac{1}{d^2}
	\end{align*}
\end{Lemma}
\begin{proof}
	Here we apply a similar method of the proof of Proposition 5 in \cite{rastegin2014notes}. Let 
	\begin{align}
	\ket{\Psi}=&\frac{1}{d^{3/2}}\sum_{j=1}^{d^2}\ket{\phi_j}\ket{\phi_j^{*}},\label{b1}
\\	\ket{\Phi_k}=&\frac{\sqrt{d+1}}{d^{3/2}}\sum_{j=1}^{d^2}\omega^{k(j-1)}\ket{\phi_j}\ket{\phi_j^{*}},\label{b2}
	\end{align}
	here $k=1,2,\cdots,d^2-1,$ and $\ket{\phi^{*}}$ is a vector such that its components are conjugate to the corresponding components of $\ket{\phi}$. The $d^2$ vectors (\ref{b1}),(\ref{b2})  construct an orthonormal basis of the space $\mathcal{H}\otimes\mathcal{H},$ hence,
	\begin{align}
	\ket{i}\bra{j}\otimes \mathbb{I}\ket{\Psi}=\sum_{k=1}^{d^2-1}a_k\ket{\Phi_k},\label{f9}
	\end{align}
	then
	\begin{align}
	a_k=&\bra{\Phi_k}\ket{i}\bra{j}\otimes \mathbb{I}\ket{\Psi}\nonumber\\
	=&\frac{\sqrt{d+1}}{d^3}\sum_{i,j=1}^{d^2}\omega^{-q(i-1)}\bra{\phi_i}\ket{i}\bra{j}\ket{\phi_j}\bra{\phi_j}\ket{\phi_i}\nonumber\\
	=&\frac{\sqrt{d+1}}{d}\sum_{i=1}^{d^2}\omega^{-q(i-1)}p_i,
	\end{align}
	here $p_i=\frac{\bra{\phi_i}\ket{i}\bra{j}\ket{\phi_i}}{d}.$ Next
	\begin{align}
	\bra{\Psi}(\ket{i}\bra{j}\otimes\mathbb{I})^2\ket{\Psi}=\frac{1}{d}\tr(\ket{i}\bra{j}i\rangle\bra{j})=0,
	\end{align}
	Through (\ref{f9}), 
	\begin{align}
	\sum_{k=1}^{d^2-1}a_k^{*}a_k=(d+1)\sum_{j=1}^{d^2}p_j^2-\frac{d+1}{d^2},
	\end{align}
	that is,
	\begin{align*}
\sum_{j=1}^{d^2}p_j^2=\frac{1}{d^2}
	\end{align*}
\end{proof}
\begin{Lemma}\label{l3}
	 Let $\{\ket{e^{\a}_i}|i=1,2,\cdots,m\}_{\a=1}^L$ be the MUBs in a $d$ dimensional system $\mathcal{H}$, and $\mathcal{N}^{(\a)}=\{Q_i^{(\a)}=\ket{e^{\a}_i}\bra{e^{\a}_i}|i=1,2,\cdots,m\}_{\a=1}^L$. Assume $\{\ket{i}|i=1,2,\cdots,d\}$ is any orthonormal base of $\mathcal{H}$, when $i\ne j,$
	\begin{align*}
	\sum_{\a}^L\sum_{i=1}^{d^2}\tr(Q_i^{(\a)}\ket{i}\bra{j})\le\frac{L}{d}
	\end{align*}
\end{Lemma}
The proof of Lemma \ref{l3} is similar to that of Lemma \ref{l2}, here we omit it. 
\end{document}